OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (3, 15)

OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (40, 5)

OpenStudy (anonymous):

y = 8x y = 5x y = 1/8x y = 1/5x

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (40, 5) A. y = 8x B. y = 5x C. y = 1/8x D. y = 1/5x